OpenStudy (anonymous):

Ramona is a saleswoman whose base pay plus commissions amounted to $52,561 last year. If her base pay was $23,960 and she made $77,300 in sales last year, what was her rate of commission? A. 37% B. 46% C. 31% D. 68%

OpenStudy (nurali):

B + C = 52561 B = 23960 C = kS S = 77300 23960 + k(77300) = 52561 k = (52561 - 23960)/77300 = 0.37 = 37%

OpenStudy (anonymous):

thank you @Nurali

OpenStudy (nurali):

Anytime.